Consider an investment of $4,000 that’s compounded continuously at a rate of 6.5%. What’s the investment worth after 11 months?
Answer #1

A= Pe^(rt)

investment of $4,000, means P =4000

a rate of 6.5%, means r= 0.065

11 months means , t =11/12 years

hence

A = 4000*e^(0.065*11/12)

=$4245.58